OpenStudy (anonymous):

An Olympic sprinter can go from a state of rest to 11 meters per second in 10 seconds. What is the average acceleration of the sprinter? A. 110 m/s² B. 110 m/s C. 1.1 m/s² D. 0.9 m/s

OpenStudy (anonymous):

it is C right

OpenStudy (anonymous):

Acceleration is (meters/sec) per second... so ... \[\frac{11 \text{m/sec}}{ 10 \text{sec}}\] is ... \[\frac{11}{10}\frac{\text{m}}{\text{sec}^2}\] ... you got it! Good job!